電気電子システム創生工学輪講及び演習 4 単位 (必修)

Electrical and Electronic Engineering Seminar and Exercise  電気電子創生工学コース教員

【授業目的】研究に関する文献や著書を読んで，研究分野に関する知識を習得する【授業概要】前期 (修士) 課程における研究テーマに関連する他の研究者による論

文を読み，所属する研究室において発表，討論する．本科目は，工業に関する科目である．

【授業形式】講義形式とポートフォーリオ形式の併用【キーワード】文献調査，発表技法，研究討論【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】特になし【履修上の注意】 [注意] 【到達目標】

1. 文献の収集法とその読み方を習得する

2. 研究討論の方法を習得する

3. 人前で自分の意見を述べる技法を習得する【授業計画】

1. 下記の 2. から 4. を 16 週繰り返す

2. テキストの選定や文献の収集

3. 文献の要約と発表

4. 発表内容に関する討論【成績評価基準】ポートフォーリオまたは試験で評価【教科書】教科書は指導教官が授業開始時に指定【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216782 【対象学生】開講コース学生のみ履修可能【連絡先】 [連絡先]

Electrical and Electronic Engineering Seminar and Exercise 4 units (compulsory)

Teacher of course

Target⟩ To get wide knowledge on your reaserach topics. Outline⟩ Students read scientific papers relating to your research and get wide

knowledge on your reaserach topics. Style⟩ Lecture in combination with Portfolio �������⟩ reading references, presentation skill, discussion Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ N/A Notice⟩ [注意]

Goal⟩

1. To get skills to collect scientific references and knowledge relating your research topic

2. To get skills on discussion about a scientific topic

3. To get presentation skills

Schedule⟩

1. Continure the followings.

2. Select text books and/or collect references.

3. Summarize the contents of a reference and/or a text book and presnet them.

4. Discuss about the contents of the presentation. Evaluation Criteria⟩ Evaluated with portfolios and/or examinations Textbook⟩ Text books and references will be specified by your supervisors at the

first class. Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216782

Student⟩ Able to be taken by only specified class(es) Contact⟩ [連絡先]